OpenStudy (anonymous):

Condense this: 9ln(x+18)-3lnx

OpenStudy (anonymous):

\[\ln((x+8)^9)-\ln(x^3)\] \[\ln(\frac{(x+8)^9}{x^3})\]

OpenStudy (anonymous):

satellite is right

OpenStudy (anonymous):

when there is ln \[a^{b}\] = b ln a and when there is ln a - ln b = ln (a/b)

OpenStudy (anonymous):

thanks